What is the integer if twice the square of a positive integer is 35 more than 9 times that integer? how do I get the integer I am lost.
Answer by ankor@dixie-net.com(12700)

What is the integer if twice the square of a positive integer is 35 more than 9 times that integer?
:
Let x = the integer
Just write what it says:
:
Twice x^2 = 35 more than than 9x
2x^2 = 9x + 35
:
Arrange as a quadratic equation:
2x^2 - 9x - 35 = 0
:
Factors to:
(2x + 5)(x - 7) = 0
:
x = 7, it said the "positive integer"
:
:
Check:
2(7^2) = 9(7) + 35
98 = 63 + 35